# Characterizing the Efficiency of Distributed Training: A Power, Performance, and Thermal Perspective

### v1

@@ Abstract.
The rapid scaling of Large Language Models (LLMs) has pushed training workloads far beyond the limits of single-node analysis, demanding a deeper understanding of how these models behave across large-scale, multi-GPU systems. In this paper, we present a comprehensive characterization of LLM training across diverse real-world workloads and hardware platforms, including NVIDIA H100/H200 and AMD MI250 GPUs. We analyze dense and sparse models under various parallelism strategies – tensor, pipeline, data, and expert – and evaluate their effects on hardware utilization, power consumption, and thermal behavior. We further evaluate the effectiveness of optimizations such as activation recomputation and compute–communication overlap. Our findings show that performance is not determined solely by scaling hardware capacity. Scale-up systems with fewer, higher-memory GPUs can outperform scale-out systems in communication-bound regimes, but only under carefully tuned configurations; in other cases, scale-out deployments achieve superior throughput. We also show that certain parallelism combinations, such as tensor with pipeline, lead to bandwidth underutilization due to inefficient data chunking, while increasing microbatch sizes beyond a certain point induces bursty execution and peak power excursions that worsen thermal throttling. These insights reveal how training performance is shaped by complex interactions between hardware, system topology, and model execution. We conclude by offering recommendations for system and hardware design to improve the scalability and reliability of future LLM systems and workloads. The source code of this project is available at https://github.com/sitar-lab/CharLLM-PPT.
@@ 1. Introduction
Large Language Models (LLMs) have driven datacenters to operate at unprecedented levels of computational demand. Models such as GPT (Brown et al., 2020), Llama (Grattafiori et al., 2024), Mixtral (Jiang et al., 2024a), Gemini (Comanici et al., 2025), and Claude (cla, ) require training across thousands of GPUs connected via high-bandwidth interconnect to maintain throughput. The shift to large-scale distributed training exposes fundamental bottlenecks in compute efficiency, communication overhead, load balancing, and energy consumption that are not well-captured by single-node analysis, necessitating a systematic characterization of these interactions.
@@
A key challenge in large-scale training is selecting an appropriate parallelism strategy. Each approach – data parallelism (DP), tensor parallelism (TP), pipeline parallelism (PP), and expert parallelism (EP) as used in Mixture-of-Experts (MoE) models – offers trade-offs in computation, memory efficiency, and communication patterns. These trade-offs are further influenced by software-level optimizations such as activation recomputation (Korthikanti et al., 2022), which reduces memory usage at the cost of additional computation, and compute-communication overlap, which improves efficiency but increases hardware resource contention (Wang et al., 2022). Moreover, training performance is highly sensitive to pipeline scheduling and microbatch size. For example, interleaved scheduling can improve utilization in PP workloads, but its effectiveness depends on network depth and synchronization barriers (Narayanan et al., 2021a). Similarly, microbatch size significantly impacts load balancing, activation memory, and synchronization overhead, in ways that vary across parallelism strategies.
@@
As the scale of AI training continues to grow, the design of modern infrastructure is becoming increasingly constrained by power consumption, thermal dynamics, and overall system efficiency (Stojkovic et al., 2024, 2025; goo, ; Lee et al., 2025a, b; Phanishayee et al., 2022; Adnan et al., 2024a). Dense GPU clusters often face thermal hotspots, power capping, and frequency throttling, reducing throughput and reliability. To address this, cloud providers invest in airflow optimization, dynamic power budgeting, and workload shaping, efforts typically decoupled from algorithmic choices like parallelism or device placement. Analyzing interactions between training techniques and system constraints is complex due to the diverse design space of models, parallelism strategies, hardware, and software.
@@
While most prior work (Rasley et al., 2020; Harlap et al., 2018; Zheng et al., 2022; Wang et al., 2024; Phanishayee et al., 2025; Tarnawski et al., 2020, 2021; Wang et al., 2023a; Jiang et al., 2024b; Liu et al., 2024) evaluates parallelism and optimizations (Dao et al., 2022; Pope et al., 2023; Wang et al., 2020; Shah et al., 2023; Won et al., 2024; Singh et al., 2025) primarily from a performance perspective, real-world deployments expose overlooked hardware constraints like thermal imbalance, throttling, and bandwidth limits. These factors, often underestimated in algorithmic or software-centric designs, can substantially degrade throughput and energy efficiency. In particular, we observe that hardware-induced variability interacts in complex ways with parallelization strategies.
@@ 2.1. Distributed Machine Learning
Infrastructure and runtime systems. The scale of modern machine learning (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2024a; cla, ) has led to the adoption of large-scale distributed training across multi-node, multi-GPU datacenters (Rasley et al., 2020; Shoeybi et al., 2019; Kuchaiev et al., 2019). These setups, such as the example illustrated in Figure 1, rely on domain-specific stacks (Paszke et al., 2017; Abadi et al., 2015; Mahajan et al., 2016; Sharma et al., 2016; Park et al., 2017), scalable frameworks such as Megatron (Shoeybi et al., 2019), DeepSpeed (Rasley et al., 2020), and NeMo (Kuchaiev et al., 2019), which introduce techniques for parallelism, memory management, and throughput optimization (Shah et al., 2023; Harlap et al., 2018; Dao et al., 2022; Rasley et al., 2020; Rajbhandari et al., 2020; Tarnawski et al., 2021). Training often spans thousands of GPUs connected via high-bandwidth interconnects like NVLink, NVSwitch (nvl, ), and InfiniBand (inf, ), stressing infrastructure across compute, communication, power, and thermal limits.
@@
Parallelism strategies. Parallelism strategies distribute models across devices with trade-offs in memory, interconnect, and utilization. Data Parallelism (DP)(Krizhevsky et al., 2017; Li et al., 2020) replicates the model but incurs gradient sync overhead. FSDP(Zhao et al., 2023) shards parameters to reduce memory and enable larger models. Tensor Parallelism (TP)(Shoeybi et al., 2019) splits computation across GPUs but demands high bandwidth. Pipeline Parallelism (PP)(Harlap et al., 2018; Huang et al., 2019; Narayanan et al., 2021b) stages the model to save memory and overlap compute, but can suffer from stalls. Expert Parallelism (EP) (Rajbhandari et al., 2022; Zhu et al., 2025; Singh et al., 2023) activates subsets of experts, introducing load imbalance and expensive all-to-all communication.
@@
Device placement. Modern LLM training often uses hybrid parallelism, combining DP, TP, PP, and EP tailored to model architecture, memory constraints, and hardware topology. Prior work on device placement and strategy search optimizes parameters like parallel widths, pipeline depth, and expert parallel configuration (Zheng et al., 2022; Jia et al., 2018; Tarnawski et al., 2021; Wang et al., 2024, 2023b), considering factors such as kernel latency, network bandwidth, and memory footprint. However, the effectiveness of these strategies can be limited by system-level constraints including thermal throttling, power capping, and interconnect bottlenecks. As LLM training pushes infrastructure limits, it becomes essential to extend parallelism strategy design to incorporate system-level considerations, including cooling efficiency, power delivery constraints, and topological layout. This highlights a critical gap: parallelism and placement strategies are rarely co-designed with awareness of physical infrastructure constraints. In this work, we bridge this gap by empirically characterizing how power, thermal behavior, and interconnect contention impact distributed LLM training, uncovering bottlenecks and challenging assumptions. This motivates the need for full-stack approaches that combine parallelism-aware scheduling with system-level modeling.
@@
Network search and communication collective optimization. Previous works leverage underlying network topology information to design logical communication patterns like all-to-all, ring, or tree-based collectives (Wang et al., 2020; Shah et al., 2023; Won et al., 2024; Cai et al., 2021), with some efforts optimizing physical topologies for various models and communication patterns, including all-reduce (Wang et al., 2023b). In contrast, our characterization emphasizes the need to co-optimize communication collectives and network topology alongside parallelism strategies and system-level factors, revealing a broader set of correlations than previously explored.
@@ 2.2. Related Work
System and runtime characterization. Prior works have explored various aspects of system characterization across hardware and software stacks. Studies on memory hierarchies (Mei and Chu, 2017) and interconnect behavior (Li et al., 2018; Jin et al., 2024) provide insights into GPU microarchitecture, while others (Wang et al., 2023c; Sahraei et al., 2023; Seemakhupt et al., 2023; Gonzalez et al., 2023; Adnan et al., 2021, 2024b; Mahajan, 2024) focus on cloud-scale systems for distributed services, serverless architectures, and remote procedure calls. Recent studies have characterized LLM training workloads in datacenters, examining scheduling, hardware failures, and learning dynamics across checkpoints (Hu et al., 2024). Pythia (Biderman et al., 2023) enables detailed analysis of LLM training dynamics through regular checkpoints and controlled replication across scales. AxoNN (Singh et al., 2025) and MegaScale (Jiang et al., 2024b) characterized performance and scalability of state-of-the-art parallelism and optimization techniques at large-scale training clusters with thousands of GPUs. While these works offer insights into large-scale training behavior, they primarily focus on software or aggregate metrics. In contrast, our work presents the first system-level characterization of distributed LLM training, analyzing performance, power, and thermal implications of parallelism strategies and optimization techniques across hardware platforms to unveil trade-offs affecting real-world efficiency.
@@ 3.1. Software Libraries and Workloads
Workloads, distribution strategies, and optimizations. We evaluate five representative LLMs (see Table 1), spanning both dense and Mixture-of-Experts (MoE) architectures, with model sizes ranging from 30B to 175B parameters. All experiments use the Pile (Gao et al., 2020) dataset, FP16 or BF16 precision, and a global batch size of 128. For LoRA finetuning and inference tasks, we use the PubMedQA (Jin et al., 2019) dataset. We vary tensor, pipeline, data, and expert parallelism configurations based on the model and hardware setup. For each model and cluster, we determine the minimal total model parallelism (Tensor Pipeline Expert) required to fit within GPU memory, then explore valid configurations, limiting tensor parallelism to within-node execution to reduce communication costs. We also evaluate Fully-Sharded Data Parallelism (FSDP) in conjunction with tensor parallelism (i.e., TP8+FSDP4), reflecting common 2D parallelism usage in large-scale training (Zhang et al., 2024; Zhao et al., 2023; AI, ; PyTorch, ). Standalone FSDP was not evaluated because of memory constraints arising from increased microbatch sizes.
@@
Mixtral-8x22B (Jiang et al., 2024a) | Mixture-of-Experts | 141B |
Mixtral-8x7B (Jiang et al., 2024a) | Mixture-of-Experts | 47B |
@@
All models, except MoEs, use distributed optimizers like ZeRO-1 (Rajbhandari et al., 2020), which partitions optimizer states across data-parallel ranks to reduce memory overhead. MoE models use standard optimizer configurations due to limitations in NeMo and Megatron-LM. We also evaluate optimization techniques, such as activation recomputation and compute-communication overlap, which affect efficiency based on model scale and system characteristics. Note that both NeMo and Megatron-LM maps parallelism in the following order: TP EP DP PP (Shoeybi et al., 2019; Nvidia, 2025a, b) This order determines how ranks are assigned and how communication occurs across devices, particularly in multi-node settings. Table 2 summarizes the parallelism and optimization configurations in our study.
@@ 3.2. Cluster Configuration
AMD Cluster. On AMD hardware, we use a ROCm-compatible Megatron-LM (Shoeybi et al., 2019) with ROCm 6.3 and RCCL 2.21 for collectives. Efficient kernels are supported through adapted forks of Apex, TransformerEngine, and FlashAttention. Power and utilization metrics are collected using AMD-SMI (Advanced Micro Devices, Inc., 2024), extended via Zeus for fine-grained profiling. The cluster consists of 4 nodes, each with 4MI250 accelerators. Each MI250 includes 2 GCDs, yielding 8 logical GPUs per node. Intra-node communication leverages xGMI, while nodes are connected via 100 Gbps InfiniBand. We include the MI250 system to study the performance and efficiency implications of chiplet-based GPU architectures, which introduce intra-package and inter-GCD communication asymmetries absent in NVIDIA’s monolithic designs. These differences enable us to examine their impact on parallelism strategies, resource utilization, and thermal behavior at scale. Due to limited availability and lower memory capacity on AMD platforms, we evaluate them using smaller versions of GPT-3 and Llama-3. Specifically, we scale down the models to fit within 4MI250 nodes, maintaining proportional relationships among key architectural parameters such as the number of layers, attention heads, and hidden dimensions. The resulting models contain approximately 30 billion parameters. Nonetheless, these runs provide valuable cross-platform insights that complement our broader analysis.
@@ 7.3. Maintaining Efficiency by Optimizing for Physical Realities of ML System
Our findings highlight a growing need to broaden the scope of optimization beyond isolated algorithmic and parallelism strategies to encompass hardware-level metrics and constraints. Hardware scaling, interconnect asymmetry, and thermal imbalance are no longer secondary effects due to the scale of modern large-scale ML training. Even in modern datacenter infrastructure with advanced cooling designs (Meta, 2024; Nvidia, 2024), temperature skew persists due to residual airflow mixing, rack constraints, and workload- induced heat concentration (Stojkovic et al., 2024, 2025; goo, ). As our analyses show, these factors can meaningfully impact the training efficiency, performance and system reliability, even in a well-balanced distributed workload such as LLMs. Conventional strategies that assume uniform hardware behavior often lead to only-partially true belief that runtime performance is also uniform. As a result, often prior works advocate one-size-fits-all optimizations—such as always overlapping compute and communication, uniformly increasing microbatch sizes, or applying aggressive pipelining. However, our analysis shows that such assumptions break down in practice. Variations in power availability, thermal profiles, and clock throttling create substantial runtime heterogeneity across devices, even for the same hardware.

### v2

@@ Abstract.
The rapid scaling of Large Language Models (LLMs) has pushed training workloads far beyond the limits of single-node analysis, demanding a deeper understanding of how these models behave across large-scale, multi-GPU systems. In this paper, we present a comprehensive characterization of LLM training across diverse real-world workloads and hardware platforms, including NVIDIA H100/H200 and AMD MI250 GPUs. We analyze dense and sparse models under various parallelism strategies – tensor, pipeline, data, and expert – and evaluate their effects on hardware utilization, power consumption, and thermal behavior. We further evaluate the effectiveness of optimizations such as activation recomputation and compute-communication overlap. Our findings show that performance is not determined solely by scaling hardware capacity. Scale-up systems with fewer, higher-memory GPUs can outperform scale-out systems in communication-bound regimes, but only under carefully tuned configurations; in other cases, scale-out deployments achieve superior throughput. We also show that certain parallelism combinations, such as tensor with pipeline, lead to bandwidth underutilization due to inefficient data chunking, while increasing microbatch sizes beyond a certain point induces bursty execution and peak power excursions that worsen thermal throttling. These insights reveal how training performance is shaped by complex interactions between hardware, system topology, and model execution. We conclude by offering recommendations for system and hardware design to improve the scalability and reliability of future LLM systems and workloads. The source code of this project is available at https://github.com/sitar-lab/CharLLM-PPT.
@@ 1. Introduction
Large Language Models (LLMs) have driven datacenters to operate at unprecedented levels of computational demand. Models such as GPT (Brown et al., 2020), Llama (Grattafiori et al., 2024), Mixtral (Jiang et al., 2024b), Gemini (Comanici et al., 2025), and Claude (Anthropic, 2024) require training across thousands of GPUs connected via high-bandwidth interconnect to maintain throughput. The shift to large-scale distributed training exposes fundamental bottlenecks in compute efficiency, communication overhead, load balancing, and energy consumption that are not well-captured by single-node analysis, necessitating a systematic characterization of these interactions.
@@
A key challenge in large-scale training is selecting an appropriate parallelism strategy. Each approach – data parallelism (DP), tensor parallelism (TP), pipeline parallelism (PP), and expert parallelism (EP) as used in Mixture-of-Experts (MoE) models – offers trade-offs in computation, memory efficiency, and communication patterns. These trade-offs are further influenced by software-level optimizations such as activation recomputation (Korthikanti et al., 2022), which reduces memory usage at the cost of additional computation, and compute-communication overlap, which improves efficiency but increases hardware resource contention (Wang et al., 2022). Moreover, training performance is highly sensitive to pipeline scheduling and microbatch size. For example, interleaved scheduling can improve utilization in PP workloads, but its effectiveness depends on network depth and synchronization barriers (Narayanan et al., 2021b). Similarly, microbatch size significantly impacts load balancing, activation memory, and synchronization overhead, in ways that vary across parallelism strategies.
@@
As the scale of AI training continues to grow, the design of modern infrastructure is becoming increasingly constrained by power consumption, thermal dynamics, and overall system efficiency (Stojkovic et al., 2024, 2025; Google, 2025; Lee et al., 2025a, b; Phanishayee et al., 2022; Adnan et al., 2024b). Dense GPU clusters often face thermal hotspots, power capping, and frequency throttling, reducing throughput and reliability. To address this, cloud providers invest in airflow optimization, dynamic power budgeting, and workload shaping, efforts typically decoupled from algorithmic choices like parallelism or device placement. Analyzing interactions between training techniques and system constraints is complex due to the diverse design space of models, parallelism strategies, hardware, and software.
@@
While most prior work (Rasley et al., 2020; Harlap et al., 2018; Zheng et al., 2022; Wang et al., 2024; Phanishayee et al., 2025; Tarnawski et al., 2020, 2021; Wang et al., 2023b; Jiang et al., 2024a; Liu et al., 2024) evaluates parallelism and optimizations (Dao et al., 2022; Pope et al., 2023; Wang et al., 2020; Shah et al., 2023; Won et al., 2024; Singh et al., 2025) primarily from a performance perspective, real-world deployments expose overlooked hardware constraints like thermal imbalance, throttling, and bandwidth limits. These factors, often underestimated in algorithmic or software-centric designs, can substantially degrade throughput and energy efficiency. In particular, we observe that hardware-induced variability interacts in complex ways with parallelization strategies.
@@ 2.1. Distributed Machine Learning
Infrastructure and runtime systems. The scale of modern machine learning (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2024b; Anthropic, 2024) has led to the adoption of large-scale distributed training across multi-node, multi-GPU datacenters (Rasley et al., 2020; Shoeybi et al., 2019). These setups, such as the example illustrated in Figure 1, rely on domain-specific stacks (Paszke et al., 2017; Abadi et al., 2015; Mahajan et al., 2016; Sharma et al., 2016; Park et al., 2017), scalable frameworks such as Megatron (Shoeybi et al., 2019), DeepSpeed (Rasley et al., 2020), and NeMo (Kuchaiev et al., 2019), which introduce techniques for parallelism, memory management, and throughput optimization (Shah et al., 2023; Harlap et al., 2018; Dao et al., 2022; Rasley et al., 2020; Rajbhandari et al., 2020; Tarnawski et al., 2021). Training often spans thousands of GPUs connected via high-bandwidth interconnects like NVLink, NVSwitch (Nvidia, 2025d), and InfiniBand (Nvidia, 2025a), stressing infrastructure across compute, communication, power, and thermal limits.
@@
Parallelism strategies. Parallelism strategies distribute models across devices with trade-offs in memory, interconnect, and utilization. Data Parallelism (DP)(Krizhevsky et al., 2017; Li et al., 2020) replicates the model but incurs gradient sync overhead. FSDP(Zhao et al., 2023) shards parameters to reduce memory and enable larger models. Tensor Parallelism (TP)(Shoeybi et al., 2019) splits computation across GPUs but demands high bandwidth. Pipeline Parallelism (PP)(Harlap et al., 2018; Huang et al., 2019; Narayanan et al., 2021a) stages the model to save memory and overlap compute, but can suffer from stalls. Expert Parallelism (EP) (Rajbhandari et al., 2022; Zhu et al., 2025; Singh et al., 2023) activates subsets of experts, introducing load imbalance and expensive all-to-all communication.
@@
Device placement. Modern LLM training often uses hybrid parallelism, combining DP, TP, PP, and EP tailored to model architecture, memory constraints, and hardware topology. Prior work on device placement and strategy search optimizes parameters like parallel widths, pipeline depth, and expert parallel configuration (Zheng et al., 2022; Jia et al., 2018; Tarnawski et al., 2021; Wang et al., 2024, 2023a), considering factors such as kernel latency, network bandwidth, and memory footprint. However, the effectiveness of these strategies can be limited by system-level constraints including thermal throttling, power capping, and interconnect bottlenecks. As LLM training pushes infrastructure limits, it becomes essential to extend parallelism strategy design to incorporate system-level considerations, including cooling efficiency, power delivery constraints, and topological layout. This highlights a critical gap: parallelism and placement strategies are rarely co-designed with awareness of physical infrastructure constraints. In this work, we bridge this gap by empirically characterizing how power, thermal behavior, and interconnect contention impact distributed LLM training, uncovering bottlenecks and challenging assumptions. This motivates the need for full-stack approaches that combine parallelism-aware scheduling with system-level modeling.
@@
Network search and communication collective optimization. Previous works leverage underlying network topology information to design communication patterns like all-to-all, ring, or tree-based collectives (Wang et al., 2020; Shah et al., 2023; Won et al., 2024; Cai et al., 2021), with some efforts optimizing physical topologies for various models and communication patterns, including all-reduce (Wang et al., 2023a). In contrast, our characterization emphasizes the need to co-optimize communication collectives and network topology alongside parallelism strategies and system-level factors, revealing a broader set of correlations than previously explored.
@@ 2.2. Related Work
System and runtime characterization. Prior works have explored various aspects of system characterization across hardware and software stacks. Studies on memory hierarchies (Mei and Chu, 2017) and interconnect behavior (Li et al., 2018; Jin et al., 2024) provide insights into GPU microarchitecture, while others (Wang et al., 2023c; Sahraei et al., 2023; Seemakhupt et al., 2023; Gonzalez et al., 2023; Adnan et al., 2021, 2024a; Mahajan, 2024) focus on cloud-scale systems for distributed services, serverless architectures, and remote procedure calls. Recent studies have characterized LLM training workloads in datacenters, examining scheduling, hardware failures, and learning dynamics across checkpoints (Hu et al., 2024). Pythia (Biderman et al., 2023) enables detailed analysis of LLM training dynamics through regular checkpoints and controlled replication across scales. AxoNN (Singh et al., 2025) and MegaScale (Jiang et al., 2024a) characterized performance and scalability of state-of-the-art parallelism and optimization techniques at large-scale training clusters with thousands of GPUs. While these works offer insights into large-scale training behavior, they primarily focus on software or aggregate metrics. In contrast, our work presents the first system-level characterization of distributed LLM training, analyzing performance, power, and thermal implications of parallelism strategies and optimization techniques across hardware platforms to unveil trade-offs affecting real-world efficiency.
@@ 3.1. Software Libraries and Workloads
Workloads, distribution strategies, and optimizations. We evaluate five representative LLMs (see Table 1), spanning both dense and Mixture-of-Experts (MoE) architectures, with model sizes ranging from 30B to 175B parameters. All experiments use the Pile (Gao et al., 2020) dataset, FP16 or BF16 precision, and a global batch size of 128. For LoRA finetuning and inference tasks, we use the PubMedQA (Jin et al., 2019) dataset. We vary tensor, pipeline, data, and expert parallelism configurations based on the model and hardware setup. For each model and cluster, we determine the minimal total model parallelism (Tensor Pipeline Expert) required to fit within GPU memory, then explore valid configurations, limiting tensor parallelism to within-node execution to reduce communication costs. We also evaluate Fully-Sharded Data Parallelism (FSDP) in conjunction with tensor parallelism (i.e., TP8+FSDP4), reflecting common 2D parallelism usage in large-scale training (Zhang et al., 2024; Zhao et al., 2023; AI, 2025; PyTorch, 2025). Standalone FSDP was not evaluated because of memory constraints arising from increased microbatch sizes.
@@
Mixtral-8x22B (Jiang et al., 2024b) | Mixture-of-Experts | 141B |
Mixtral-8x7B (Jiang et al., 2024b) | Mixture-of-Experts | 47B |
@@
All models, except MoEs, use distributed optimizers like ZeRO-1 (Rajbhandari et al., 2020), which partitions optimizer states across data-parallel ranks to reduce memory overhead. MoE models use standard optimizer configurations due to limitations in NeMo and Megatron-LM. We also evaluate optimization techniques, such as activation recomputation and compute-communication overlap, which affect efficiency based on model scale and system characteristics. Note that both NeMo and Megatron-LM maps parallelism in the following order: TP EP DP PP (Shoeybi et al., 2019; Nvidia, 2025b, c) This order determines how ranks are assigned and how communication occurs across devices, particularly in multi-node settings. Table 2 summarizes the parallelism and optimization configurations in our study.
@@ 3.2. Cluster Configuration
AMD Cluster. On AMD hardware, we use a ROCm-compatible Megatron-LM (Shoeybi et al., 2019) with ROCm 6.3 and RCCL 2.21 for collectives. Efficient kernels are supported through adapted forks of Apex, TransformerEngine, and FlashAttention. Power and utilization metrics are collected using AMD-SMI (Advanced Micro Devices, Inc., 2024), extended via Zeus for fine-grained profiling. The cluster consists of 4 nodes, each with 4MI250 accelerators. Each MI250 includes 2 GCDs, yielding 8 logical GPUs per node. Intra-node communication leverages xGMI, while nodes are connected via 100 Gbps InfiniBand. We include the MI250 system to study performance and efficiency implications of chiplet-based GPU architectures, which introduce intra-package and inter-GCD communication asymmetries absent in NVIDIA’s monolithic designs. These differences enable us to examine their impact on parallelism strategies, resource utilization, and thermal behavior at scale. Due to limited availability and lower memory capacity on AMD platforms, we evaluate them using smaller versions of GPT-3 and Llama-3. Specifically, we scale down the models to fit within 4MI250 nodes, maintaining proportional relationships among key architectural parameters such as the number of layers, attention heads, and hidden dimensions. The resulting models contain approximately 30 billion parameters. Nonetheless, these runs provide valuable cross-platform insights that complement our broader analysis.
@@ 7.3. Maintaining Efficiency by Optimizing for Physical Realities of ML System
Our findings highlight a growing need to broaden the scope of optimization beyond isolated algorithmic and parallelism strategies to encompass hardware-level metrics and constraints. Hardware scaling, interconnect asymmetry, and thermal imbalance are no longer secondary effects due to the scale of modern large-scale ML training. Even in modern datacenter infrastructure with advanced cooling designs (Meta, 2024; Nvidia, 2024), temperature skew persists due to residual airflow mixing, rack constraints, and workload- induced heat concentration (Stojkovic et al., 2024, 2025; Google, 2025). As our analyses show, these factors can meaningfully impact the training efficiency, performance and system reliability, even in a well-balanced distributed workload such as LLMs. Conventional strategies that assume uniform hardware behavior often lead to only-partially true belief that runtime performance is also uniform. As a result, often prior works advocate one-size-fits-all optimizations—such as always overlapping compute and communication, uniformly increasing microbatch sizes, or applying aggressive pipelining. However, our analysis shows that such assumptions break down in practice. Variations in power availability, thermal profiles, and clock throttling create substantial runtime heterogeneity across devices, even for the same hardware.